OpenStudy (anonymous):

1/5 ÷ 4/5 = ??? 1/4 4/25 5/9 3/10

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }×\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

answer will need to be simplified

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

that is it simplified

OpenStudy (anonymous):

Heres a tip if they already have the same denominator just do it normally with the numerators

OpenStudy (anonymous):

1|20

OpenStudy (anonymous):

dont give answers please it's against the coc

OpenStudy (anonymous):

@iceicebaby lease dont give answers help them

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

5×2? 5×4?

OpenStudy (anonymous):

1\4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/5 ÷ 2/7 = ??? 7/10 1 3/7 1 2/35

OpenStudy (anonymous):

and thats suppose to be 5×1 not 2

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }×\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

1|10

OpenStudy (anonymous):

7|10

OpenStudy (anonymous):

5/7 ÷ 4/5 = ??? 5/14 25/28 1 3/25 9/11

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

place

OpenStudy (anonymous):

5/9 ÷ 4/5 = ??? 9/13 25/36 5/18 1 11/25

OpenStudy (anonymous):

1.52 ÷ 2 0.76 0.86 1.14 1.01

OpenStudy (anonymous):

wold that be 0.76